# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ Acknowledgements
We would like to thank Avik Chakraborty and Mrityunjay Nath for helpful discussions on various stages of this work. PD acknowledges the financial support by DST-INSPIRE fellowship (IF200253), Department of Science and Technology, Govt. of India. The work of MM is supported by DST-FIST initiative (SR/FST/LS-I/2020/621) and (GrantSR/FST/PSI-225/2016).
@@ Appendix A Bit threads: a brief review
Here we will briefly review the notion of bit threads proposed by Headrick and Freedman [3]. We will start from the continuation version of well known max flow-min cut theorem originated in network theory.
@@
where, $\mathrm{C}$ is a positive quantity. A ‘surface’ is defined as a (d1)(d-1)-dimensional oriented submanifold. Then, flux of the flow through such a surface ‘$m$’ can be expressed as
@@
The inequalities (A.4) provide an an upper bound for the flux of any flow via the area of a minimal surface $m(A)$ (homologous to $A$), which is the bottleneck (as shown in fig. 2) [^13] We thank Ayushi Suman for the figure.. The MFMC theorem (A.6) implies that $n_{a}\upsilon^{a}=\mathrm{C}$, and hence $\upsilon^{a}=\mathrm{C}n^{a}$, everywhere on $m(A)$.
@@ Appendix B Covariant phase space formalism
The covariant phase space formalism [50, 15, 16] is an approach to study Hamiltonian mechanics in which one constructs a phase space by taking the set of solutions of the equation of motions instead of the generalised coordinates and their conjugates. To study the entropy and thermodynamics of black holes, people mainly use the Iyer-Wald formalism [15]. There have been many modifications and generalisations to this formalism. One such example is [24], where the authors considered the boundary Lagrangian separately. However, the entropy obtained from both of these formalisms is same. Here, we will briefly discuss both of these formalisms, the Wald entropy and the well known first law of black hole thermodynamics.
@@ B.1 Iyer-Wald Formalism
In CPS, the phase space is considered to be a manifold $\mathcal{M}$ with a symplectic two-form $\Omega$. Then from the symplectic form, one can define the Hamiltonian. The action for a generally invariant theory is
@@
where, $\boldsymbol{\mathrm{L}}$ is the Lagrangian d-form for any d-dimensional theory, and $\Phi$ denotes all the dynamical fields collectively. Taking variation of this action leads to
@@
where $\delta\Phi$ represents a generic field perturbation and provides the basis of (tangent space of) the phase space $\mathcal{P}$. The equations of motion are given by the set of equations $\boldsymbol{\mathrm{E}}_{\Phi}=0$. The quantity $\mathbf{\Theta}_{\rm{IW}}$ is called the symplectic potential. It is a (d1)(d-1)-form in spacetime $\mathcal{M}$ and one form on the phase space, i.e., $\mathbf{\Theta}_{\rm{IW}}$ is a (d1;1)(d-1;1)-form. Then one can define the quantity called (pre)symplectic current as
@@
Here, $\delta_{1}\Phi$ and $\delta_{2}\Phi$ are two arbitrary field perturbations, $\mathrm{d}$’ denotes the exterior derivative over the spacetime $\mathcal{M}$ and $\delta$’ represents the exterior derivative over the phase space manifold $\mathcal{P}$. Using this notation, it is evident that $\boldsymbol{\omega}_{\rm{IW}}$ is a two-form on the phase space, or more compactly, is a (d1;2)(d-1;2)-form. A symplectic form then can be constructed by integrating the (pre) symplectic current over a codimension-one spacelike surface $\Sigma$,
@@
Now, given a transformation generated by $\epsilon$, one can define different quantities like Noether charge, Noether current, and the Hamiltonian (or, more generally a conserved charge). For diffeomorphism invariant theories, covariance tells that $\delta_{\xi}\Phi=\mathcal{L}_{\xi}\Phi$, where $\xi$ is the diffeomorphism generator. So the (pre)-symplectic current becomes,
@@
The above expression can be written in terms of the Noether current and Noether charge as well. Now, we will talk about the recent modified version of the covariant phase space with boundaries.
@@ B.2 CPS with Boundaries
Recently Harlow and Wu generalised the CPS formalism by adding a boundary term and the result is the covariant phase space formalism with boundaries (CPSB) [24]. In this kind of formulation, the action is written as,
@@
with boundary conditions for $\Phi^{i}(x)$ on the spatial boundary $\Gamma$. Here, $\Phi$ is dynamical and $\Xi$ is background field. Then one imposes that the action is well posed, i.e., $\Phi$ obeys EOM iff $\mathrm{S}$ is stationary upto future/past terms at $\Phi$. Similar to IW formalism, here also the Lagrangian is considered to be a $d$-form (for $d$-dimensional spacetime/manifold) instead of a scalar. Therefore, the variation of the Lagrangian can be written as,
@@
where, $\mathrm{d}$’ is exterior derivative on space-time, $\mathbf{\Theta}_{\rm{HW}}$ is called symplectic potential. Thus, the variation of the action is given by
$\delta\mathrm{S}=\int_{\mathcal{M}}\boldsymbol{\mathrm{E}}_{i}\delta\Phi^{i}+\int_{\Sigma_{\pm}}\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}+\int_{\Gamma}\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}$ (B.11)
As mentioned earlier, the action is required to be stationary. The stationarity of $\mathrm{S}$ implies $\delta\mathrm{S}=0$. This condition results to two important equations:
@@
and, the other one is the non-trivial one,
$\big{(}\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big{)}\mid_{\Gamma}=\mathrm{d}\boldsymbol{\mathrm{C}}$ (B.13)
for any (d2)(d-2)-form $\boldsymbol{\mathrm{C}}$. Now, one can define a quantity named pre-symplectic current, $\boldsymbol{\omega}_{\rm{HW}}\equiv\delta(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathrm{C}})\mid_{\tilde{\mathcal{P}}}$ with certain properties. Integrating the pre-symplectic current over a Cauchy slice one obtains the pre-symplectic form,
@@
where, $X_{\xi}\equiv\int d^{d}\mathrm{x}\;\mathcal{L}_{\xi}\;\Phi^{i}(x)\dfrac{\delta}{\delta\Phi^{i}(x)}$. Noether current is defined as,
@@
$X_{\xi}.\tilde{\mathbf{\Omega}}_{\rm{HW}}=\delta\left(\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big{(}\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\big{)}\right).$ (B.17)
@@
$H_{\xi}=\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big{(}\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\big{)}.$ (B.18)
This is true for any kind of higher derivative theory. Some examples can be seen in [24]. There are some special theories (like the theory of general relativity), where $\boldsymbol{\mathrm{L}}$ is covariant under arbitrary $\xi^{\mu}$. This gives us $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$ and $\boldsymbol{\mathrm{J}}_{\xi}$ becomes an exact form. So, one can define a (d2)(d-2)-form $\boldsymbol{\mathrm{Q}}_{\xi}$ such that,
@@
$H_{\xi}=\int_{\partial\Sigma}\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathrm{C}}\right),$ (B.20)
which shows that the Hamiltonian is a totally boundary term.
@@ B.3 Black Hole Thermodynamics
where, $\kappa$ is the surface gravity of the black hole. The obvious relation between these new generators is given by,
@@
Now, one can define the conserved charges corresponding to the generators. The conserved charges for time evolution and rotation are defined by integrating the codimensional-2 currents $\mathbf{k}_{\eta_{M}}$ and $\mathbf{k}_{\eta_{J_{i}}}$ at the spatial infinity,
@@
The Cauchy surface $\Sigma$ has boundary $\partial\Sigma=\mathcal{S}_{\infty}\cup(-\mathcal{B})$, where $\mathcal{B}$ is a surface (here, the bifurcation surface) homologous to $\mathcal{S}_{\infty}$ (the spatial infinity). As $\xi_{H}$ is a Killing vector, any variation due to this should vanish, i.e., $\mathcal{L}_{\xi_{H}}\Phi=0$, or more preciously,
@@
The temperature of the horizon is defined as $T_{\rm{H}}=\frac{\kappa}{2\pi}$. Substituting it, one obtains the well known 1st law of black hole thermodynamics,
@@
Appendix C Flow in D11-brane background
$D1$-brane is an object in $10D$ supergravity. An effective theory for this system can be obtained in 3 dimensions by compactifying it on a 7-sphere. This effective $3D$ Einstein-dilaton system is governed by the action [51],
$\mathrm{S}_{\rm{ED}}=\dfrac{1}{16\pi G_{N}}\int d^{3}x\sqrt{-g}\left[R-\dfrac{\nu}{2}\partial_{\mu}\varphi\partial^{\mu}\varphi-{\cal{V}}(\varphi)\right],$ (C.1)
where, $\nu=\frac{16}{9}$, $\varphi$ is the dilaton and its potential ${\cal{V}}(\varphi)=-\frac{24}{L^{2}}e^{\frac{4}{3}\varphi}$. The $3D$ metric of the $D1$-brane is given by,
@@ Appendix C Flow in D11-brane background
$\varphi=-3\ln{\left(\dfrac{r}{L}\right)}.$ (C.4)
The metric (C.2) contains two constants: the AdS radius $L$ and the non-extremal parameter $r_{0}$. We can use either of these two or both to do the parametric variations as discussed in section 2. For simplicity we will use the AdS radius $L$ for the variatio. Unlike previous examples, this solution consists of both the gravity and a scalar (dilaton) sector. For the gravity part we can use (2.20) to find the non-zero component of the codim-2 current:
@@
where the suffix $EH$ denotes the Einstein-Hilbert part of the action. For the scalar sector we can use the expression [23],
$k^{\alpha\beta}_{Sc}=\dfrac{1}{8\pi G_{N}}\left[\xi^{\beta}\tilde{\nu}\nabla^{\alpha}\phi\delta\phi-(\alpha\leftrightarrow\beta)\right],$ (C.6)
@@
By adding these two different contributions we get the total codim-2 current,
@@
Now, to check if it gives rise to bit threads, we need to see whether this vector is proportional to the unit normal of a RT surface corresponding to a boundary subregion. Let us first try to find the RT surface equation. Finding this surface for an interval (in the boundary) is akin to going away from the centre of the bulk, i.e., taking the limit $r\rightarrow\infty$ (or $r\gg r_{0}$). In this limit the metric (C.2) reduces to
@@
We can try to further simplify it. Substituting $y=r^{4}$ yields
@@
RT surface can be obtained by integrating over the following differential equation,
@@
where, $y_{0}$ is the deepest point of the RT surface in the bulk. Integrating both sides of the above equation gives us the RT surface equation
@@
[…]
@@
This vector field also does not matches with the unit normal (C.15) at the RT surface. So, we need to add a non-trivial vector field $v^{a}_{\mathrm{d}}$ or the harmonic function $\varphi$. Solving the Laplace’s equation in the background (C.12) and using proper boundary conditions, we find
$\varphi=\dfrac{1}{4G_{N}}\left(x-\dfrac{1}{2\sqrt{y_{0}}}E\left(\dfrac{1}{2}\sec^{-1}(\dfrac{y}{y_{0}})\,\bigg{|}\,\frac{1}{p_{x}^{2}}\right)-\dfrac{\mathfrak{K}}{32}\sqrt{y}-\dfrac{4y_{0}y^{3/2}}{(y^{2}-y_{0}^{2})^{3/2}}\left(\dfrac{x^{2}}{2}-x\mathfrak{f}(y)\right)\right),$ (C.17)
with $\mathfrak{f}(y)=\frac{\sqrt{y^{2}-y_{0}^{2}}}{2y_{0}\sqrt{y}}{{}_{2}}F_{1}\left(\,\frac{1}{4}\,,\,1\,;\,\frac{3}{4}\,;\,\frac{y^{2}}{y_{0}^{2}}\,\right)$.

### v2

@@ Acknowledgments
We would like to thank Debajyoti Sarkar, Avik Chakraborty, and Mrityunjay Nath for helpful discussions on various stages of this work. We are also grateful to the anonymous referee for his/her constructive review and helpful comments. PKD acknowledges the financial support from the DST-INSPIRE fellowship (IF200253), Department of Science and Technology, Govt. of India. This work is also supported by DST-FIST initiative (SR/FST/LS-I/2020/621) and (GrantSR/FST/PSI-225/2016).
@@ Appendix A Bit threads: a brief review
Here we will briefly review the notion of bit threads proposed by Headrick and Freedman [27]. We will start from the continuation version of the well-known max-flow min-cut theorem, which originated in network theory.
@@
where, $\mathrm{C}$ is a positive quantity. A ‘surface’ is defined as a (d1)(d-1)-dimensional oriented submanifold. Then, the flux of the flow through such a surface ‘$m$’ can be expressed as
@@
The inequalities (A.4) provide an upper bound for the flux of any flow via the area of a minimal surface $m(A)$ (homologous to $A$), which is the bottleneck (as shown in figure 2). [^17] We thank Ayushi Suman for the figure. The MFMC theorem (A.6) implies that $n_{a}\upsilon^{a}=\mathrm{C}$, and hence $\upsilon^{a}=\mathrm{C}n^{a}$, everywhere on $m(A)$.
@@ Appendix B Covariant phase space formalism
The covariant phase space formalism [19, 39, 58] is an approach to study Hamiltonian mechanics in which one constructs a phase space by taking the set of solutions to the equations of motion instead of the generalised coordinates and their conjugates. To study the entropy and thermodynamics of black holes, people mainly use the Iyer-Wald formalism [39]. There have been many modifications and generalisations to this formalism. One such example is [30], where the authors considered the boundary Lagrangian separately. However, the entropy obtained from both of these formalisms is the same. Here, we will briefly discuss both of these formalisms, the Wald entropy and the well-known first law of black hole thermodynamics.
@@ B.1 Iyer-Wald Formalism
[…]
@@
where, $\boldsymbol{\mathrm{L}}$ is the Lagrangian $d$-form for any $d$-dimensional theory, and $\Phi$ denotes all the dynamical fields collectively. Taking a variation of this action leads to
@@
where $\delta\Phi$ represents a generic field perturbation and provides the basis of (tangent space of) the phase space $\mathcal{P}$. The equations of motion are given by the set of equations $\boldsymbol{\mathrm{E}}_{\Phi}=0$. The quantity $\mathbf{\Theta}_{\rm{IW}}$ is called the symplectic potential. It is a (d1)(d-1)-form in spacetime $\mathcal{M}$ and one form on the phase space, i.e., $\mathbf{\Theta}_{\rm{IW}}$ is a (d1;1)(d-1;1)-form. Then, one can define the quantity called (pre)symplectic current as
@@
Here, $\delta_{1}\Phi$ and $\delta_{2}\Phi$ are two arbitrary field perturbations, $\mathrm{d}$’ denotes the exterior derivative over the spacetime $\mathcal{M}$ and $\delta$’ represents the exterior derivative over the phase space manifold $\mathcal{P}$. Using this notation, it is evident that $\boldsymbol{\omega}_{\rm{IW}}$ is a $2$-form on the phase space, or more compactly, is a (d1;2)(d-1;2)-form. A symplectic form then can be constructed by integrating the symplectic current over a codimension-one spacelike surface $\Sigma$,
@@
Now, given a transformation generated by $\xi$, one can define different quantities like Noether charge, Noether current, and the Hamiltonian (or, more generally, a conserved charge). For diffeomorphism invariant theories, covariance tells that $\delta_{\xi}\Phi=\mathcal{L}_{\xi}\Phi$, where $\xi$ is the diffeomorphism generator. So the (pre)-symplectic current becomes,
@@
The above expression can be written in terms of the Noether current and Noether charge as well. Now, we will talk about the recently modified version of the covariant phase space with boundaries.
@@ B.2 CPS with Boundaries
Recently, Harlow and Wu generalised the CPS formalism by adding a boundary term, and the result is the covariant phase space formalism with boundaries (CPSB) [30]. In this kind of formulation, the action is written as,
@@
with boundary conditions for $\Phi^{i}(x)$ on the spatial boundary $\Gamma$. Here, $\Phi$ is dynamical and $\Xi$ is the background field. Then one imposes that the action is well posed, i.e. $\Phi$ obeys EOM iff $\mathrm{S}$ is stationary (up to some boundary terms in future and past slices). Similar to the IW formalism, here also the Lagrangian is considered to be a $d$-form (for $d$-dimensional spacetime manifold) instead of a scalar. Therefore, the variation of the Lagrangian can be written as
@@
where, $\mathrm{d}$’ is the exterior derivative in space-time, $\mathbf{\Theta}_{\rm{HW}}$ is called symplectic potential. Thus, the variation of the action is given by
$\delta\mathrm{S}=\int_{\mathcal{M}}\boldsymbol{\mathrm{E}}_{i}\delta\Phi^{i}+\int_{\Sigma_{\pm}}\big(\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big)+\int_{\Gamma}\big(\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big),$ (B.11)
where, the boundary of the manifold $\mathcal{M}$ is: $\partial\mathcal{M}=\Gamma\cup\Sigma_{-}\cup\Sigma_{+}$. $\Sigma_{\pm}$ are the future (past) boundaries and $\Gamma$ is the spatial boundary. As mentioned earlier, the action must be stationary. The stationarity of $\mathrm{S}$ implies $\delta\mathrm{S}=0$. This condition results in two important equations:
@@
and the other one is the non-trivial one,
$\big(\mathbf{\Theta}_{\rm{HW}}+\delta\boldsymbol{\ell}\big)\mid_{\Gamma}=\mathrm{d}\boldsymbol{\mathsf{C}}$ (B.13)
for any (d2)(d-2)-form $\boldsymbol{\mathsf{C}}$. Now, one can define a quantity named pre-symplectic current, $\boldsymbol{\omega}_{\rm{HW}}\equiv\delta(\mathbf{\Theta}_{\rm{HW}}-\mathrm{d}\boldsymbol{\mathsf{C}})\mid_{\tilde{\mathcal{P}}}$ with certain properties. Integrating the pre-symplectic current over a Cauchy slice, one obtains the pre-symplectic form,
@@
where, $X_{\xi}\equiv\int d^{d}\mathrm{x}\;\mathcal{L}_{\xi}\;\Phi^{i}(x)\dfrac{\delta}{\delta\Phi^{i}(x)}$. The Noether current is defined as,
@@
$X_{\xi}.\tilde{\mathbf{\Omega}}_{\rm{HW}}=\delta\left(\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big(\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathsf{C}}\big)\right).$ (B.17)
@@
$H_{\xi}=\int_{\Sigma}\boldsymbol{\mathrm{J}}_{\xi}+\int_{\partial\Sigma}\big(\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathsf{C}}\big).$ (B.18)
This is true for any kind of higher derivative theory. Some examples can be found in [30]. There are some special theories (like the theory of general relativity), where $\boldsymbol{\mathrm{L}}$ is covariant under arbitrary $\xi^{\mu}$. This gives us $\mathrm{d}\boldsymbol{\mathrm{J}}_{\xi}=0$ and $\boldsymbol{\mathrm{J}}_{\xi}$ becomes an exact form. So, one can define a (d2)(d-2)-form $\boldsymbol{\mathrm{Q}}_{\xi}$ such that
@@
$H_{\xi}=\int_{\partial\Sigma}\left(\boldsymbol{\mathrm{Q}}_{\xi}+\xi.\boldsymbol{{\ell}}-X_{\xi}.\boldsymbol{\mathsf{C}}\right),$ (B.20)
which shows that the Hamiltonian is a total boundary term.
@@ B.3 Black Hole Thermodynamics
where $\kappa$ is the surface gravity of the black hole. The obvious relation between these new generators is given by,
@@
Now, one can define the conserved charges corresponding to the generators. The conserved charges for time evolution and rotation are defined by integrating the codim-$2$ currents $\mathbf{k}_{\eta_{M}}$ and $\mathbf{k}_{\eta_{J}}$ at the spatial infinity,
@@
The Cauchy surface $\Sigma$ has boundary $\partial\Sigma=\mathcal{S}_{\infty}\cup(-\mathcal{B})$, where $\mathcal{B}$ is a surface (here, the bifurcation surface) homologous to $\mathcal{S}_{\infty}$ (the spatial infinity). As $\xi_{H}$ is a Killing vector, any variation along its flow should vanish, i.e., $\mathcal{L}_{\xi_{H}}\Phi=0$, or more precisely,
@@
The temperature of the horizon is defined as $T_{\rm{H}}=\frac{\kappa}{2\pi}$. Substituting it, one obtains the well-known 1st law of black hole thermodynamics,
@@
Appendix C Flow in D1-brane background
$D1$-brane is an object in $10D$ Type IIB supergravity. An effective theory for this system can be obtained in 3 dimensions by compactifying it on a 7-sphere. This effective $3D$ Einstein-dilaton system is governed by the action [21],
$\mathrm{S}_{\rm{ED}}=\dfrac{1}{16\pi G_{N}}\int d^{3}x\sqrt{-g}\left[R-\dfrac{\nu}{2}\partial_{\mu}\vartheta\partial^{\mu}\vartheta-{\cal{V}}(\vartheta)\right],$ (C.1)
where, $\nu=\frac{16}{9}$; $\vartheta$ is the dilaton and its potential ${\cal{V}}(\vartheta)=-\frac{24}{L^{2}}e^{\frac{4}{3}\vartheta}$. The $3D$ metric of the $D1$-brane is given by,
@@ Appendix C Flow in D1-brane background
$\vartheta=-3\ln{\left(\dfrac{r}{L}\right)}.$ (C.4)
The metric (C.2) contains two constants: the AdS radius $L$ and the non-extremal parameter $r_{0}$. We can use either of these two or both to do the parametric variations as discussed in section 2. For simplicity, we will use the AdS radius $L$ as the variation parameter. Unlike previous examples, this solution comprises both the gravitational and a scalar (dilaton) sector. For the gravity part, we can use (2.20) to find the non-zero component of the codim-2 current:
@@
where the suffix $EH$ denotes the Einstein-Hilbert part of the action. For the scalar sector, we can use the expression [29],
$k^{\alpha\beta}_{Sc}=\dfrac{1}{8\pi G_{N}}\left[\xi^{\beta}\tilde{\nu}\nabla^{\alpha}\vartheta\delta\vartheta-(\alpha\leftrightarrow\beta)\right],$ (C.6)
@@
By adding these two different contributions, we get the total codim-2 current,
@@
Now, to check if it gives rise to bit threads, we need to see whether this vector is proportional to the unit normal of an RT surface corresponding to a boundary subregion. Let us first try to find the RT surface equation. Finding this surface for an interval (in the boundary) is akin to going away from the centre of the bulk, i.e., taking the limit $r\rightarrow\infty$ (or $r\gg r_{0}$). In this limit the metric (C.2) reduces to
@@
We can try to simplify it further. Substituting $y=r^{4}$ with $L=1$ yields
@@
The RT surface can be obtained by integrating over the following differential equation,
@@
where $y_{0}$ is the deepest point of the RT surface in the bulk. Integrating both sides of the above equation gives us the RT surface equation
@@
Clearly, the vector (C.10) does not match the above unit normal. One reason is that we evaluated the vector field in the original 3-dimensional metric (C.2), whereas the unit normal or the RT surface itself is obtained in the limit $r>>r_{0}$. So, we need to find the vector in this limit. Proceeding the same way as before, we find the vector field obtained from CPS as
@@
This vector field also does not match the unit normal (C.15) at the RT surface. So, we need to add a non-trivial vector field $v^{a}_{\mathrm{d}}$ or the harmonic function $\varphi$. Solving Laplace’s equation in the background (C.12) and using proper boundary conditions, we find
$\varphi=\dfrac{1}{4G_{N}}\left(x-\dfrac{1}{2\sqrt{y_{0}}}E\left(\dfrac{1}{2}\sec^{-1}\left(\dfrac{y}{y_{0}}\right)\,\bigg|\,\frac{1}{p_{x}^{2}}\right)-\dfrac{\mathfrak{K}}{32}\sqrt{y}-\dfrac{4y_{0}y^{3/2}}{(y^{2}-y_{0}^{2})^{3/2}}\left(\dfrac{x^{2}}{2}-x\mathfrak{f}(y)\right)\right),$ (C.17)
with $\mathfrak{f}(y)=\frac{\sqrt{y^{2}-y_{0}^{2}}}{2y_{0}\sqrt{y}}{{}_{2}}F_{1}\left(\,\frac{1}{4}\,,\,1\,;\,\frac{3}{4}\,;\,\frac{y^{2}}{y_{0}^{2}}\,\right)$ and $E(\mathsf{k}|\mathsf{m})$ is the incomplete elliptic integral of second kind.